Calculate the pH of 0.34 M ZnCl2 solution.

Express your answer using two decimal places.
Answer #1

ZnCl2 is a salt and it dissociate as:

ZnCl2 --> Zn2+ + 2Cl-

In aqueous solution we have

Zn2+ + 6H2O ----> [Zn(H2O)6]2+

It dissociate as:

[Zn(H2O)6]2+ <=> [Zn(H2O)5OH]+ + H+

From literature, Ka for [Zn(H2O)6]2+ is 2.5 *10-10

[Zn(H2O)6]2+ <=> [Zn(H2O)5OH]+ H+
I(M) 0.34 0 0
C(M) -x +x +x
E(M) 0.34-x x x

Ka = [H+][Zn(H2O)5OH]+ / [Zn(H2O)6]2+

2.5 *10-10 = x2 /(0.34-x)

x2 + (2.5 *10-10 )x - (8.5 *10-11) = 0

x = 9.22 *10-6 = [H+]

pH = -log[H+] = - log (9.22 *10-6 ) = 5.035